### v3

# GraphSnapShot: Graph Machine Learning Acceleration with Fast Storage and Retrieval

Abstract
Graph learning has become essential across various domains, yet current sampling techniques face significant challenges when scaling to dynamic or large-scale networks. Popular methods, such as GraphSAGE and Cluster-GCN, encounter issues like excessive memory usage, neighbor explosion, and limited adaptability to dynamic changes. To address these limitations, we introduce GraphSnapShot, a framework specifically designed for efficient and scalable graph learning through dynamic snapshot caching.
GraphSnapShot incorporates the GraphSDSampler, a key module that dynamically captures, updates, and stores local graph snapshots, optimizing both memory usage and computational efficiency. The framework offers innovative caching strategies, including Fully Cache Refresh (FCR), On-The-Fly (OTF), and Shared Cache Mode, tailored to balance computational demands and cache freshness. OTF methods dynamically update only portions of the graph cache, combining static pre-sampling with real-time dynamic resampling, significantly improving computational speed without compromising accuracy.
We validate the effectiveness of GraphSnapShot on well-known graph learning datasets, including OGBN-arxiv, OGBN-products, OGBN-mag, as well as Citeseer, Cora, and Pubmed. Experimental results show that GraphSnapShot achieves significant reductions in GPU memory usage and computational time compared to traditional samplers, such as those in DGL. Notably, GraphSnapShot maintains competitive accuracy while efficiently handling complex multi-hop neighborhoods in large-scale graphs.
This framework is particularly effective for applications such as social network analysis, bioinformatics, and recommendation systems. Its ability to process evolving graph structures makes it a valuable tool for researchers and practitioners. The code for GraphSnapShot is publicly available at code link111https://github.com/NoakLiu/GraphSnapShot.
1University of Wisconsin-Madison
2University of Notre Dame
1 Introduction
In our recent research, we have developed a framework called GraphSnapShot, which has been proven an useful tool for graph learning acceleration. The core idea of GraphSnapShot is to capture and update the state of local graph structures dynamically, just like taking snapshots of graphs.
GraphSnapShot is designed to efficiently capture, store and update the dynamic snapshots of graph data, enabling us to track patterns in the structure of graph networks. This technique is useful for most graph learning tasks that relies on topology analysis or networks are constantly evolving, such as social media analysis, biological networks, or any system where the relationships between entities change over time.
The key components of GraphSnapShot is the GraphSDSampler. GraphSDSampler can efficiently capture, update, retrieve and store graph snapshots of topology while doing computation at the same time, which makes graph learning computation significantly faster.
In experiments, GraphSnapShot shows efficiency. It can promote computation speed significantly compared to traditional NeighborhodSamplers implemented in dgl, and it can reduce the GPU & memory usage during training, with little loss of accuracy. Experimental results show that the GraphSnapShot has potential to be a powerful tool for large graph training acceleration.
2 Background
Local structure analysis in graph learning applications are fundamental for understanding and analyzing complex networks in various domains. This concept pertains to the idea of capturing the localized patterns and relationships within a large graph structure. Efficient graph structure learning are crucial in various applications, from social network analysis to bioinformatics, where useful hidden pattern are discovered by graph structure learning.
The GraphSnapShot framework we proposed provide insight to study graph learning by using dynamically updated local graph snapshots. GraphSnapShot is particularly valuable in large-scale networks where global analysis can be computationally intensive and less informative for certain types of inquiries. By focusing on local structures, researchers can detect community patterns, identify influential nodes, and understand local clustering behaviors in a more efficient way.
Applications of local structure learning are diverse and impact several fields:
Social Network Analysis: Understanding individual or entity interactions within a network to reveal social dynamics and community formation.
Bioinformatics: In biological networks, local structures assist in identifying functional modules or predicting protein interactions.
Recommendation Systems: Improving recommendation accuracy by focusing on immediate user-item interaction patterns.
Network Security: Analyzing local patterns for detecting anomalies or potential security breaches.
3 Motivation
In machine learning, masks are powerful tools for reducing computation and enhancing model efficiency. In attention masks of long sequences, models selectively attend to relevant tokens within sequences, bypassing unnecessary computations and enabling more focused processing. Masks can also find utility in regularization techniques like dropout, where they aid in randomly deactivating units during training, thus preventing overfitting and promoting generalization. Additionally, in tasks such as instance segmentation in computer vision, masks efficiently delineate object boundaries, streamlining the computational load by restricting processing to specific regions of interest. Through these mechanisms, masks can reduce computational resources and contribute to the scalability of machine learning models. Notably, masked language modeling (Song et al., 2019; Wettig et al., 2023) and masked vision learning (Li et al., 2023) showcase the significance of masks.
Although masks are widely used in NLP and computer vision, their applications in graph learning are limited. In this study, we propose GraphSnapShot, a novel method for cache-masked graph structure learning. The aim of GraphSnapShot is to reduce computational resources and accelerate the speed of graph machine learning at large scale.
4 Model Construction
GraphSnapShot introduces caching masked graph structure for graph learning acceleration. The main process of GraphSnapShot involves two phases: preprocess sampling and dynamic resampling, both are methods for caching graph mask.
Traditional graph learning methods struggle with extensive memory usage and slow computation. Node-wise sampling methods, including well-known ones like GraphSAGE (Hamilton et al., 2018), are not optimized for multi-hop scenarios due to the computational resources usage that grows exponentially with each additional hop— a phenomenon known as the "neighbor explosion." This makes them unsuitable for large-scale graphs where the capture of extended neighborhoods is crucial. Layer-wise sampling, such as in FastGCN(Chen et al., 2018), although designed to mitigate this explosion by sampling nodes per layer, often loses valuable connectivity information, leading to a sparse and inaccurate representation of the graph’s multi-hop structure. Furthermore, subgraph sampling methods such as Cluster-GCN (Chiang et al., 2019) focus on localized computations and overlook vital inter-subgraph connections, failing to capture the broader topology necessary for a comprehensive multi-hop domain analysis.
GraphSnapShot’s core idea is to achieve efficient cache management in graph learning, it is designed to maintain an up-to-date cache through the following strategies:
Fully Cache Refresh (FCR): Periodically refresh the whole cache to get a new representative of the graph local structure.
On-The-Fly (OTF): Retreive the new representation of graph local structure in each computation, with partially retrieve from the previous graph local structure and partially retrieve from the disk. In detail the OTF can be identified as follows based on their fetch and refresh methods:
Partial Cache Refresh and Full Fetch: Dynamically updates portions of the cache or retrieves all data from cache for computation as needed.
Partial Refresh with Partial Fetch: OTF method in partially updating and partially retrieving data from both disk and cache, balancing cache up-to-date cache with computational efficiency.
Only Fetch Mode: Fetch data for computation from cache by OTF mode with update existing cache periodically.
Only Refresh Mode: Updates the existing cache by OTF mode.
Shared Cache Mode: Unlike traditional methods that cache individual nodes or edges, this mode involves caching significant portions of the graph that are frequently accessed by various processes. This strategy optimizes memory usage and ensures that the most frequently accessed data is readily available, enhancing the responsiveness and efficiency of graph machine learning tasks.
GraphSnapShot is designed to solve the inefficacy for current graph masking methods on large graphs. Typical distributed graph processing System such as Marius (Mohoney et al., 2021), will re-sample all the local structure each time, and load corresponding node embeddings from disk, which is time-consuming. GraphSnapShot proposing a new method for quick local structure retrieval by integrating the benefits of both static and dynamic processing. In the preprocessing phase, a static graph snapshot is sampled and stored in memory from the disk. Then, in the computation phase, the algorithm re-samples the graph snapshot, using a mix of data partially from memory and partially from disk. Concurrently, a smaller, dynamic snapshot is sampled from the disk. This dynamic snapshot is crucial as it is used to swap with the portion of the graph stored in memory. This process ensures the computation can always learn "fresh" snapshot without overwhelming the memory capacity. GraphSnapShot allows the algorithm to continuously update and adapt to changes in the graph while maintaining a manageable memory footprint, making it an effective tool for processing large, dynamic graph data.
In the GraphSnapShot system, disks are used to store edge and node embedding information, while computational resources (CPU/GPU) are utilized for neural network computations (such as backpropagation) and local structure masking. Additionally, the cache is employed to store the K-V pairs information of the sampled structure.
Disk: Persistent storage of graph structure information, such as edges and node embedding parameters. Those embeddings are stored long-term and loaded into memory when needed for further processing and analysis.
Cache: Store frequent K-V pairs of GraphSnapShot to accelerate graph learning. These K-V pairs include the identifiers of the sampled structure nodes and their embedding parameters, allowing for more efficient repeated access to the graph data.
CPU/GPU: Computational Resources to responsible for performing local structure computations on the graph, including neural network forward and backward propagation as well as the dynamic resampling and masking of the graph’s local structure.
5 Algorithmic Design
As for the cache retrieval and refreshing methods, we develop three methods to simulate different strategies.
5.1 FBL
The first method is FBL, which is short for full batch load. FBL is the traditional method where most graph processing systems current are deployed. Those system does not caching any information and resample graph structure each time from disk to memory for computation.
5.2 OTF
The "On The Fly" (OTF) method, updates portions of the cache dynamically while leaving other segments of the cache information unchanged. In detailed implementation of OTF, we design 4 different mode of it for real world application. They are (Full Fetch, Partial Fetch)x(Full Refresh, Partial Refresh).
In *OTF-refresh* mode, we implement a partial cache refresh strategy, which dynamically updates only a fraction of the cache based on a defined refresh rate.
5.2.1 Algorithmic Design for OTF-refresh
Let denote a heterogeneous graph, where and represent the vertices and edges, respectively. The vertices are categorized by types, specified by hete_label.
: An array indicating the desired number of neighbors to sample at each depth level.
: A scaling factor for fanouts to determine the initial cache size for each depth.
: The interval after which the cache is considered for partial refreshing.
: Refresh rate defining the fraction of the cache to update during each refresh cycle.
The OTF refresh strategy enhances the algorithm’s responsiveness to changes in the graph’s structure by updating a subset of the cache, thus balancing between computational efficiency and sampling accuracy. This approach is particularly beneficial in dynamic environments where frequent full cache refreshes are impractical.
In *OTF-fetch* mode, we design an algorithm that can maintaining up-to-date neighborhood information in dynamic graph environments. Traditional methods that rely solely on static pre-sampling can lead to stale data, affecting the model’s performance. OTF-fetch mode introduces a fetching mechanism that balances between pre-sampled cache and real-time graph data, enhancing accuracy and efficiency.
5.2.2 Algorithmic Design for OTF-fetch
Let be a heterogeneous graph, where and represent the vertices and edges, respectively. Each vertex belongs to a specific type, categorized under hete_label.
: An array indicating the number of neighbors to sample at each layer.
amp_rate: Amplification rate that determines the size of the initial cache.
: The period after which the cache is partially fetched to adjust the sample data dynamically.
fetch_rate: Determines the fraction of the cache to be dynamically fetched alongside real-time data from the graph.
The OTF fetching strategy introduces a significant enhancement to graph sampling by allowing dynamic data retrieval, which integrates cached and real-time data. This mechanism ensures that the graph samples are both current and computationally efficient to retrieve, catering particularly well to environments where the graph’s structure may change frequently. Moreover, this approach reduces the overhead associated with frequent full cache refreshes, thereby optimizing both memory usage and processing time.
5.3 Algorithmic Design for OTF-partial fetch & partial refresh
Let represent a heterogeneous graph where and denote the sets of vertices and edges, respectively, with vertices categorized by types specified by hete_label.
: Specifies the number of neighbors to sample at each depth.
amp_rate: Amplification rate determining the initial cache size for each layer.
: Period after which the cache is refreshed.
refresh_rate: Proportion of the cache to be refreshed dynamically during each cycle.
The OTF partial fetch and refresh strategy significantly enhances the responsiveness and accuracy of graph sampling in dynamic environments. By selectively updating the cache, the algorithm reduces computational overhead while maintaining high data freshness, crucial for downstream applications like machine learning where timely and accurate data is vital.
5.4 OTF partial refresh & full fetch
OTF partial refresh & full fetch is a mechanism for both partial cache refresh and full fetch. This methodology optimizes the sampling process by dynamically adjusting cache content to reflect the latest graph state, thus balancing computational efficiency with data freshness.
Efficient graph sampling is pivotal in applications involving dynamic graphs, especially in scenarios where graph structures evolve over time. Traditional static caching methods often lead to outdated data, adversely affecting the performance of graph-based algorithms. Our approach, through dynamic cache management, ensures timely updates of the sampled data, significantly enhancing the relevance of the data used in downstream processes.
Let denote a heterogeneous graph where and represent vertices and edges, respectively, categorized by hete_label.
: An array specifying the number of neighbors to sample at each depth.
amp_rate: Amplification rate that determines the initial size of the cache for each depth.
: Period after which the cache undergoes a refresh operation.
refresh_rate: Proportion of the cache that is refreshed in each cycle.
5.5 FCR
The third method is FCR, which is short for "fully cache refreshing" in a batch. For FCR, in each batch, the the structure information are sampled from cache for computation, and between each batch, the cache is fully refreshed.
The Fully Cache Refresh (FCR) strategy effectively scales the depth and breadth of pre-sampled neighborhoods stored in the cache. This mechanism reduces the need for frequent neighborhood sampling from the graph, thereby lowering computational overhead and enhancing performance, especially in dynamic and large-scale heterogeneous graph environments. The periodic refresh rate ensures the cache remains up-to-date and reflects recent changes in the graph structure.
Let be a heterogeneous graph where and represent vertices and edges respectively. The vertices may belong to different types, specified by hete_label.
: An array indicating the number of neighbors to sample at each depth of the sampling process.
: Fractional caching rate, controlling the scaling of fanouts for pre-sampling.
: The interval after which the cache is refreshed to incorporate recent graph changes.
Traditional methods often suffer from high computational costs due to repetitive neighborhood sampling. Our approach mitigates these issues by integrating a Fully Cache Refresh, which dictates the extent of neighborhood pre-sampling and cache utilization.
The Fully Cache Refresh approach, while ensuring the freshness of the sampled data, can be computationally intensive compared to partial or incremental refresh strategies, as it involves re-sampling large portions of the graph frequently. Therefore, it’s crucial to carefully configure the refresh period T to balance performance with computational overhead, particularly in very large or rapidly-changing graphs.
This approach is ideal for scenarios where the accuracy and timeliness of the graph data are critical, such as in real-time monitoring systems or dynamic social networks where relationships and attributes may frequently change. By refreshing the entire cache periodically, the FCR method ensures that the sampling process remains robust and relevant to the current graph structure, thus supporting high-quality graph-based analyses and applications.
5.6 Shared Cache Design
Graph sampling is often computationally intensive, especially when dealing with large-scale heterogeneous graphs. The shared cache design aims to mitigate these challenges by maintaining a cache of pre-sampled neighborhoods, thus facilitating rapid access and reducing the frequency of expensive graph traversal operations.
Let denote a heterogeneous graph where and represent the sets of vertices and edges, respectively. fanouts defines the number of neighbors to sample at each depth of sampling, and represents the scaling factor for cache size determination.
: Specifies the number of neighbors to sample at each layer.
: Amplification factor applied to each fanout to determine the initial cache size.
: Period after which the cache is refreshed to ensure it remains representative of the graph’s evolving structure.
The shared cache design effectively reduces the computational overhead associated with frequent and intensive graph sampling operations. By leveraging an amplified fanout strategy, the cache maintains a ready pool of neighborhood data that can be rapidly accessed during the sampling process, significantly enhancing performance and efficiency. The periodic refresh mechanism ensures the cache’s data remains current, accommodating changes in the graph’s topology over time.
6 GraphSnapShot Overview
Our approach involves splitting a graph into two subgraphs based on the node degree threshold. Nodes with a degree higher than a specified threshold are categorized into a dense subgraph, while those with a degree below the threshold form a sparse subgraph. This division allows for the application of different sampling techniques to each subgraph, optimizing the computational resources and processing time.
For dense subgraphs, we utilize advanced sampling methods in GraphSnapShot such as FCR, OTF, FCR-SC, OTF-SC, etc, which are designed to efficiently manage the memory cache while maintaining the quality of the sampled neighborhood. These samplers dynamically adjust their sampling rates and cache refresh rates based on the learning phase and node centrality, ensuring efficient processing of dense regions.
For sparse subgraphs, we employ a Full Batch Load (FBL) approach, akin to DGL’s NeighborSampler. This method is suitable for areas of the graph with lower degrees, where the computational overhead is inherently less, allowing for direct processing without the need for intricate sampling techniques.
We use the Deep Graph Library (DGL(Wang et al., 2020)) framework to implement new samplers in GraphSnapShot, leveraging its efficient handling of graph data structures and operations. Our implementation involves loading the graph, applying the graph splitting based on node degree, and then processing each subgraph with the respective sampling technique. We evaluated the performance of our new models on datasets in ogbn-benchmark(Hu et al., 2021). The results show that GraphSnapShot can significantly reduce training time and memory usage without compromising the quality of GNN training.
Bridging the Gap Between Pure Dynamic Algorithms and Static Memory Storage with GraphSnapShot:
Dynamic graph algorithms, such as GraphSAGE, which require resampling the entire graph for each computation, incur significant overhead. This is primarily due to the need for constant, full-graph resampling to accommodate the ever-changing nature of dynamic graphs. GraphSnapShot addresses this issue by establishing a static snapshot at preprocessing phase and use dynamic processing to update partial snapshot at computation. This approach significantly reduces the computational burden by providing a stable, memory-stored snapshot (sG) as a baseline. GraphSnapShot ensures efficient computation and maintains the flexibility needed to handle dynamic changes in the graph.
Tradeoff Between Dynamic Sampling and Resampling:
In GraphSnapShot, the balance between dynamic sampling (DynamicSample) and resampling (ReSample) plays a pivotal role in maintaining an up-to-date representation of the graph. Dynamic sampling involves integrating new nodes from the disk into memory, capturing the latest updates and changes. Resampling, on the other hand, focuses on sampling and processing existing nodes from memory, which are then rotated back to disk storage. However, if DynamicSample is too large, it is close to traditional sampling, resulting in prolonged disk retrieval times. Conversely, while resampling processes memory-stored nodes, an excessive ReSample may overly rely on memory data, potentially leading to a decrease in accuracy. Striking this balance is essential: moderating dynamic sampling to prevent inefficiencies and ensuring resampling maintains graph accuracy without an undue dependence on memory. In short, this is a tradeoff between optimizing memory usage (quick accessing) and model accuracy.
Local Snapshot Swap Strategy:
The Add and Remove functions are designed to simulate the swap of nodes between memory and disk, reflecting the dynamic shifts between active and less active parts of the graph. This local snapshot swapping strategy is crucial for handling large-scale dynamic graphs, as it reduces the demand on memory and computational resources.
In conclusion, the design of our GraphSnapShot Process algorithm considers the complexity and variability of dynamic graph analysis. By combining static sampling with dynamic resampling, the algorithm effectively manages and analyzes graphs whose structures evolve over time, demonstrating robust performance and flexibility in handling large-scale network analyses.
Dense & Sparse Graphs Processing:
Our approach shows significant improvements in computational efficiency compared to traditional methods. The dual sampling strategy for dense and sparse graphs allows for targeted resource allocation, effectively handling dense and sparse regions within the graph. By using FBL, sparse graph processing can avoid resampled on same node set; by using GraphSnapShot, dense graph processing can have smaller cache size and quicker processing time. This results in faster training times without compromising the accuracy of the GNN models. Belows are figures to visulize the sample efficiency for Dense & Sparse Processing.
7 Experimental Result
7.1 Theoretial Analysis of Traditional Disk-Memory Strategy and Our Disk-Cache-Memory Strategy
In the traditional Graph System like Marius(Mohoney et al., 2021), a disk-cache model is deployed, where the graph structure is resampled each time from the disk into the memory for computation. In our current GraphSnapShot model, we deploy a disk-cache-memory model where K-V pair for reused graph structure are stored in the cache to promote the system performance.
Theoretically, if we denote the batch processing size as , the cache size as and the cache refreshing rate as , and the processing speed of cache as , and the processing speed of memory as . Then we can get the batch average processing time for disk-memory model as and the batch average processing time for disk-cache-memory model as .
7.2 Training Time Reduction Analysis
The provided table presents a detailed comparison of training time acceleration for various computational methods relative to the baseline method FBL. It quantitatively illustrates the percentage reduction in training times when employing FCR, FCR-shared cache, OTF, and OTF-shared cache methods across three distinct settings: [20, 20, 20], [10, 10, 10], and [5, 5, 5]. Each entry reflects the efficiency gains achieved by the respective method, highlighting how advancements in computational techniques can significantly decrease the time required for model training.
| Method/Setting | [20, 20, 20] | [10, 10, 10] | [5, 5, 5] |
| FCR | 7.05% | 14.48% | 13.76% |
| FCR-shared cache | 7.69% | 14.33% | 14.76% |
| OTF | 11.07% | 23.96% | 23.28% |
| OTF-shared cache | 13.49% | 25.23% | 29.63% |
The training time comparison table reveals that the OTF-shared cache method consistently provides the most substantial time accelerations, achieving up to 29.63% faster training times than the FBL method, particularly in the smallest setting. This performance underscores its effectiveness in scenarios demanding quick model training. The FCR methods, both with and without shared caching, offer steady improvements, enhancing training speeds by approximately 7% to 14.76%. Notably, the OTF method shines in smaller settings, suggesting increased efficiency in less complex computational tasks. The positive impact of shared caching across the FCR and OTF methods indicates that optimized resource management can lead to significant reductions in training times, making these methods highly beneficial in high-performance computing environments where reducing time is essential.
7.3 Runtime Memory reduction Analysis
From the visualizations and data provided in the table of runtime memory reduction analysis, it is evident that the computational methods FCR, FCR-shared cache, OTF, and OTF-shared cache demonstrate significant improvements in reducing runtime memory usage compared to the baseline FBL method up to 90%.
7.4 GPU Reduction Analysis
This significant reduction in GPU memory usage demonstrates the effectiveness of the GraphSnapShot methodology. By intelligently balancing the trade-off between full storage of sparse graph regions and sampling of dense graph regions, the proposed approach greatly enhances memory efficiency, which is particularly beneficial for large-scale graph processing tasks.
| Dataset | Original | Optimized | Compression |
| ogbn-arxiv | 1,166,243 | 552,228 | 52.65% |
| ogbn-products | 123,718,280 | 20,449,813 | 83.47% |
| ogbn-mag | 5,416,271 | 556,904 | 89.72% |
7.5 Fluent Generation Analysis
The partial fetch and partial refresh mode are designed to achieve fluent generation in GraphSnapShot. As in the GraphSnapShot, the most important part of the way is to develop a great opportunity to achieve …
| Method | Description | Time (seconds) | |
| Average | Worst | ||
| Full Refresh (FR) + OTF Partial Fetch (PF) | Most Common | - | - |
| Partial Fetch (PF) + Partial Refresh (PR) | Less Common | 0.23 | 0.82 |
| Partial Refresh (PR) + Partial Fetch (PF) | Less Common | 0.54 | 0.54 |
| Partial Refresh (PR) + Full Fetch (FF) | Less Common | 0.31 | 0.31 |
Full Refresh (FR) + OTF Partial Fetch (PF) is the most commonly used method, leveraging full refresh alongside on-the-fly partial fetch, likely offering advantages in performance and efficiency, although no time data is provided. Partial Fetch (PF) + Partial Refresh (PR) exhibits a shorter average time of 0.23 seconds but a longer worst time of 0.82 seconds, indicating good performance in most cases but potential delays in certain scenarios due to its implementation of partial fetch and partial refresh. Partial Refresh (PR) + Partial Fetch (PF) has an average time of 0.54 seconds and a worst time of 0.54 seconds, showing more consistent performance across various scenarios, being closest to a graph structure query with each refresh. Partial Refresh (PR) + Full Fetch (FF) shows both an average and worst time of 0.31 seconds, similar to the second method in average time but excelling in worst-case scenarios, while also offering the best fluent generation due to its implementation of partial refresh and full fetch.
7.6 Experimental Analysis of SSDReS Performance
In our experiments, we explored the efficacy of the GraphSnapShot algorithm by testing the tradeoff between static and dynamic resampling. Static fetch (proportional to ) focuses on quick access to graph topology information stored in memory. Dynamic resampling (proportional to ), on the other hand, is a measurement for the memory swap, it addresses the real-time changes in the graph, frequently requiring access to disk-stored data and swap in the memory to update the in-memory topology.
Analyzing the outcomes of these experiments provides insights into the algorithm’s performance:
Accuracy and fetch rate: Our results showed a general increase in accuracy as fetch rate increased, which shows dynamic resampling has critical roles in enhancing model precision. A higher rate of dynamic resampling implies a greater reliance on the stable structure of the graph, aiding in capturing patterns in graph topology. A relative high dynamic resampling rate is beneficial for the model’s training stability and generalizability.
Loss and fetch rate: Loss decreased with an increase in fetch rate, further validating the importance of dynamic resampling in improving model performance. Lower loss indicates reduced error in data fitting, suggesting that static resampling helps the algorithm to learn and predict the graph’s structure more accurately. This could also imply that dynamic resampling offers a stable learning environment, potentially reducing the risk of overfitting.
Training Time and refresh rate: Our experiments indicated the existence of an optimal refresh rate value between 0 and 1, which minimizes training time, and with a little loss in acuracy. We call this point as "tradeoff" point between static and dynamic resampling. Over-dependence on dynamic resampling (low refresh rate) could increase the computational burden due to frequent updates required for rapidly changing graph structures, while excessive reliance on static resampling (high refresh rate) might cause the model to overly learn from limited portion of the graph structure in memory, deduce its ability to effectively learn from the entire graph structure.
The effectiveness of the GraphSnapShot algorithm lies in the way that it deploys disk-cache-memory architecture to store and update intermediate graph local structure (GraphSnapShot). By caching the K-V pairs of the reused graph structure, the graph machine learning is significantly accelerated.
The GraphSnapShot excels in processing dynamic graph data and tasks that require exploration of graph structures, particularly for applications with complex and large-scale network structures.
Appendix A Appendix
A.1 dgl Experiments
A.1.1 datasets
| Feature | ARXIV | PRODUCTS | MAG |
| Type | Citation Net. | Product Net. | Acad. Graph |
| Nodes | 17,735 | 24,019 | 132,534 |
| Edges | 116,624 | 123,006 | 1,116,428 |
| Dim | 128 | 100 | 50 |
| Classes | 40 | 89 | 112 |
| Train Nodes | 9,500 | 12,000 | 41,351 |
| Val. Nodes | 3,500 | 2,000 | 10,000 |
| Test Nodes | 4,735 | 10,019 | 80,183 |
| Task | Node Class. | Node Class. | Node Class. |
A.1.2 Training Time Acceleration
This table provides a comparison of the training times (“training time”) of different computational methods (FBL, FCR, FCR-shared cache, OTF, OTF-shared cache) across three settings ([20, 20, 20], [10, 10, 10], and [5, 5, 5]). It also calculates the time acceleration percentage relative to FBL for methods other than FBL, expressed as a percentage.
| Method | Setting | Time (s) | Acceleration (%) |
| FBL | [20, 20, 20] | 0.2766 | - |
| [10, 10, 10] | 0.0747 | - | |
| [5, 5, 5] | 0.0189 | - | |
| FCR | [20, 20, 20] | 0.2571 | 7.05 |
| [10, 10, 10] | 0.0639 | 14.48 | |
| [5, 5, 5] | 0.0163 | 13.76 | |
| FCR-shared cache | [20, 20, 20] | 0.2554 | 7.69 |
| [10, 10, 10] | 0.0640 | 14.33 | |
| [5, 5, 5] | 0.0161 | 14.76 | |
| OTF | [20, 20, 20] | 0.2460 | 11.07 |
| [10, 10, 10] | 0.0568 | 23.96 | |
| [5, 5, 5] | 0.0145 | 23.28 | |
| OTF-shared cache | [20, 20, 20] | 0.2393 | 13.49 |
| [10, 10, 10] | 0.0559 | 25.23 | |
| [5, 5, 5] | 0.0133 | 29.63 |
| Method | Detailed Settings |
| FCR | [20, 20, 20], alpha=2, T=50 |
| [10, 10, 10], alpha=2, T=50 | |
| [5, 5, 5], alpha=2, T=50 | |
| FCR-shared cache | [20, 20, 20], alpha=2, T=50 |
| [10, 10, 10], alpha=2, T=50 | |
| [5, 5, 5], alpha=2, T=50 | |
| OTF | [20, 20, 20], amp_rate=2, refresh_rate=0.15, T=50 |
| [10, 10, 10], amp_rate=2, refresh_rate=0.15, T=50 | |
| [5, 5, 5], amp_rate=2, refresh_rate=0.15, T=50 | |
| OTF-shared cache | [20, 20, 20], amp_rate=2, refresh_rate=0.15, T=50 |
| [10, 10, 10], amp_rate=2, refresh_rate=0.15, T=50 | |
| [5, 5, 5], amp_rate=2, refresh_rate=0.15, T=50 |
OTF-shared cache shows the highest time acceleration at the smallest setting ([5, 5, 5]), with a 29.63% improvement, highlighting its efficiency for smaller datasets.
Both FCR and FCR-shared cache methods consistently perform better than FBL, especially in medium and smaller settings, indicating significant improvements in training time.
The time acceleration percentages are notably higher in the smallest setting for all methods, suggesting that these optimized approaches are particularly effective for smaller datasets.
A.1.3 Runtime Memory Reduction
| Method | Setting | RM (MB) | Reduction (%) |
| FBL | [20]*3 | 6.33 | 0.00% |
| [10]*3 | 4.70 | 0.00% | |
| [5]*3 | 4.59 | 0.00% | |
| FCR | [20]*3, ar=2, T=50 | 2.69 | 57.46% |
| [10]*3, ar=2, T=50 | 2.11 | 55.04% | |
| [5]*3, ar=2, T=50 | 1.29 | 71.89% | |
| FCR-SC | [20]*3, ar=2, T=50 | 4.42 | 30.13% |
| [10]*3, ar=2, T=50 | 2.62 | 44.15% | |
| [5]*3, ar=2, T=50 | 1.66 | 63.79% | |
| OTF | [20]*3, ar=2, rr=0.15, T=358 | 4.13 | 34.80% |
| [20]*3, ar=2, rr=0.15, T=50 | 4.19 | 33.86% | |
| [10]*3, ar=2, rr=0.15, T=50 | 1.87 | 60.07% | |
| [5]*3, ar=2, rr=0.15, T=50 | 0.32 | 93.02% | |
| OTF-SC | [20]*3, ar=2, rr=0.15, T=50 | 1.41 | 77.68% |
| [10]*3, ar=2, rr=0.15, T=50 | 0.86 | 81.58% | |
| [5]*3, ar=2, rr=0.15, T=50 | 0.67 | 85.29% |
The reduction percentage is calculated by comparing the runtime memory usage of a method under a specific setting to the baseline method’s usage at the same setting. The formula used to compute the reduction is given by:
where:
Base RM represents the runtime memory (in MB) of the baseline method for a given setting, which is the FBL method in this case.
Current RM represents the runtime memory of the current method under the same setting.
Baseline Setting (FBL): All calculations are compared against the FBL method as the baseline, where the reduction is always 0% because it is compared against itself.
FCR Method:
For the setting [20, 20, 20], memory reduction is 57.46%, showing significant effectiveness of this method at larger settings.
For the settings [10, 10, 10] and [5, 5, 5], memory reductions are 55.04% and 71.89% respectively, indicating effective memory reduction across various settings, especially at smaller scales.
FCR-shared cache Method:
This method typically uses more memory than FCR in a shared cache situation but still shows significant memory reductions, especially in the setting [5, 5, 5], with a reduction of 63.79%.
OTF Method:
For the setting [20, 20, 20], whether at a longer cycle (T=358) or a shorter cycle (T=50), memory reductions are around 34%, indicating that the change in response time does not significantly affect memory reduction.
At smaller settings ([5, 5, 5]), memory is reduced by up to 93.02%, showing extremely high efficiency of the OTF method in handling small-scale tasks.
OTF-shared cache Method:
This method shows very high memory reduction rates across all settings, particularly in [5, 5, 5] with a reduction of 85.29%, and 77.68% in [20, 20, 20], indicating that shared caching significantly enhances memory efficiency in the OTF method.
A.1.4 GPU Usage Reduction
| Threshold | Sparse Graph | Resampled Graph | Total Storage |
| 10 | 86,053 | 948,495 | 1,034,548 |
| 20 | 223,031 | 431,520 | 654,551 |
| 30 | 333,599 | 223,305 | 556,904 |
| 40 | 423,513 | 128,715 | 552,228 |
| 50 | 491,462 | 85,125 | 576,587 |
| 60 | 543,682 | 62,415 | 606,097 |
| 70 | 585,752 | 48,810 | 634,562 |
| 80 | 619,468 | 39,855 | 659,323 |
| 90 | 649,094 | 33,255 | 682,349 |
| 100 | 674,965 | 28,290 | 703,255 |
| 110 | 698,646 | 24,225 | 722,871 |
| 120 | 720,183 | 20,940 | 741,123 |
| 130 | 738,536 | 18,420 | 756,956 |
| 140 | 755,044 | 16,335 | 771,379 |
| Threshold | Sparse Graph | Resampled Graph | Total Storage |
| 10 | 92,784 | 30,942,345 | 31,035,129 |
| 20 | 491,202 | 26,278,605 | 26,769,807 |
| 30 | 1,122,314 | 23,034,195 | 24,156,509 |
| 40 | 1,894,456 | 20,644,980 | 22,539,436 |
| 50 | 2,781,440 | 18,723,690 | 21,505,130 |
| 60 | 3,797,306 | 17,072,610 | 20,869,916 |
| 70 | 4,943,892 | 15,588,690 | 20,532,582 |
| 80 | 6,230,698 | 14,219,115 | 20,449,813 |
| 90 | 7,697,986 | 12,920,310 | 20,618,296 |
| 100 | 9,327,842 | 11,708,340 | 21,036,182 |
| 110 | 11,063,622 | 10,622,880 | 21,686,502 |
| 120 | 12,882,860 | 9,671,610 | 22,554,470 |
| 130 | 14,737,002 | 8,839,755 | 23,576,757 |
| 140 | 16,667,254 | 8,088,615 | 24,755,869 |
| Threshold | Sparse Graph | Resampled Graph | Total Storage |
| 10 | 432,912 | 4,934,910 | 5,367,822 |
| 20 | 1,335,788 | 2,321,400 | 3,657,188 |
| 30 | 2,059,531 | 1,307,445 | 3,366,976 |
| 40 | 2,610,841 | 813,090 | 3,423,931 |
| 50 | 3,055,313 | 533,295 | 3,588,608 |
| 60 | 3,422,953 | 363,120 | 3,786,073 |
| 70 | 3,717,725 | 257,385 | 3,975,110 |
| 80 | 3,962,156 | 187,365 | 4,149,521 |
| 90 | 4,159,659 | 140,670 | 4,300,329 |
| 100 | 4,322,855 | 107,835 | 4,430,690 |
A.2 PyTorch Experiments
| Dataset | Nodes | Edges | Features | Classes |
| PubMed | 19,717 | 44,338 | 500 | 3 |
| Cora | 2,708 | 5,429 | 1,433 | 7 |
| CiteSeer | 3,312 | 4,732 | 3,703 | 6 |
| Operation | Duration (s) | Simulation Frequency |
| Simulated Disk Read | 5.0011 | 0.05 |
| Simulated Disk Write | 1.0045 | 0.05 |
| Simulated Cache Access | 0.0146 | 0.05 |
| In-Memory Computation | Real Computation | Real Computation |
| Operation | k_h_sampling | k_hop_retrieval | k_h_resampling |
| Disk Read | |||
| Disk Write | |||
| Memory Access |
| Dataset | Alpha | Presampled | Resampled | Sampled Depth |
| CiteSeer | 0.1, …, 0.9 | 100 | 40 | 1, 2, 3, 4 |
| Cora | 0.1, …, 0.9 | 100 | 40 | 1, 2, 3, 4 |
| PubMed | 0.1, …, 0.9 | 100 | 40 | 1, 2, 3, 4 |
| Dataset | Alpha | Presampled | Resampled | Sampled Depth |
| CiteSeer | 0.1, …, 0.9 | 20 | 10 | 1, 2, 3, 4 |
| Cora | 0.1, …, 0.9 | 20 | 10 | 1, 2, 3, 4 |
| PubMed | 0.1, …, 0.9 | 20 | 10 | 1, 2, 3, 4 |
| Abbreviation | Description |
| Adjust | Adjusts read and write costs based on system load. |
| Estimate | Estimates query cost based on read and write operations. |
| Optimize | Optimizes query based on context (’high_load’ or ’low_cost’). |
| Modify Load | Modifies query for high load optimization. |
| Modify Cost | Modifies query for cost efficiency optimization. |
| Log | Logs an I/O operation for analysis. |
| Get Log | Returns the log of I/O operations. |
| Method | Description |
| init | Initialize the buffer manager with capacity. |
| load | Load data into the buffer. |
| get | Retrieve data from the buffer. |
| store | Store data in the buffer. |